Find S[5] for the arithmetic sequence with
a[3]=39 and a[5] =61

Answer by stanbon(75887)

Find S[5] for the arithmetic sequence with
a[3]=39 and a[5] =61
----
a(3) = a + 2d = 39
a(5) = a + 4d = 61
----
Subtract to solve for "d":
2d = 22
d = 11
-------
Solve for "a":
a + 2d = 39
a = 39-2*11
a = 17
--------
To find S(5)
S(n) = (n/2)(a+L)
S(5) = (5/2)(17+61) = 5(39) = 195
=========================
Cheers,
Stan H.